How many three digit numbers can be made using only the digits 1 to 9 so that at least two of the digits are the same? For example, 112 and 737 are two such numbers.
(A) 100 (B) 200 (C) 216 (D) 225 (E) None of these

Answer by stanbon(48558)

How many three digit numbers can be made using only the digits 1 to 9 so that at least two of the digits are the same? For example, 112 and 737 are two such numbers.
---
# with >= 2 the same = (total # of 3 digit numbers) - (number with no repeats)
---
# of 3 digit with no restrictions = 9^3
# of 3 digit with no repeats = 9*8*7 =
-----
So, # with >= 2 the same = 9^3-9*8*7 = 9(9^2-56)= 9(25) = 225
=========================
Cheers,
Stan H.
=========================
(A) 100 (B) 200 (C) 216 (D) 225 (E) None of these